Every day of the week, Aaron walks to his school which is 6/10 of a kilometre away from his house. On Friday, he is 5/6 of the way to school when he realises that he forgot his Programming textbook at home and turns back to get it. Once Aaron picks up his textbook from home, he walks straight to school. How far will Aaron have walked in total that morning when he reaches the school? Write the answer in km.
1
Expert's answer
2019-05-06T14:55:44-0400

Firstly, find amount km which Aaron realises that he forgot textbook: 5/6*6/10=1/2 km. He must walk the same distance to return to the home: additional 1/2km. And the last part is length of the way to school: 6/10 km. Therefore the total distance which Aaron walked is: 1/2+1/2+6/10=16/10=1.6 km.